OpenStudy (anonymous):

Can someone explain to me how to solve this? Find a cofunction with the same value as the given expression. 18) cos 57° 19) tan 27° 20) csc 42°

OpenStudy (anonymous):

\[\cos u =\sin[ (\frac{ \pi }{ 2 })-u]\]

OpenStudy (anonymous):

:( im sorry im still confused

OpenStudy (anonymous):

cos u represents cos 57 question 18. It is equal to the cofunction on the right hand side

OpenStudy (anonymous):

im sorry i still dont get it